Lamar drove to the mountains last weekend. There was heavy traffic on the way there, and the trip took
8
hours. When Lamar drove home, there was no traffic and the trip only took
5
hours. If his average rate was
21
miles per hour faster on the trip home, how far away does Lamar live from the mountains?

Answer by stanbon(75887)

Lamar drove to the mountains last weekend. There was heavy traffic on the way there, and the trip took 8 hours.
When Lamar drove home, there was no traffic and the trip only took 5 hours.
If his average rate was 21 miles per hour faster on the trip home, how far away does Lamar live from the mountains?
----------------------
From Home::
time = 8 hrs ; rate = x mph ; distance = 8x miles
-------
To Home::
time = 5 hrs ; rate = x+21 mph ; distance = 5(x+21) = 5x + 105 miles
--------
Equation:
8x = 5x + 105
3x = 105
x = 35 mph
-----
Ans: distance to mountains = 8x = 8*35 = 280 miles
-------------------
Cheers,
Stan H.
-------------------